Session: Oct 2007 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Michael Lindqvist at Thu Nov 1 07:09:19 2007) Minor failures :
Mark5 had to be restarted. Lost one scan. Wind+rain last hours. Otherwise no known problems.

Westerbork (Tony Foley at Thu Nov 1 12:26:55 2007) Minor failures : TSYS
Single dish was used. Please change the coordinates in the vexfile accordingly. Some Tsys values are crazy; it looks like the noise source did not work. However when I tested it by hand it worked OK. Gain of BBC01 unstable. Otherwise OK.

Torun (Kaz Borkowski at Fri Nov 2 07:26:09 2007) Success :
No known problems.

Jodrell_Bank (Alastair Gunn at Fri Nov 2 11:55:51 2007) Success :
The MkII telescope was used (Jb2). No known problems.

Noto (P. Cassaro at Sun Nov 4 00:10:21 2007) Success :
No known problems

Hartebeesthoek (Jonathan Quick at Mon Nov 5 07:34:08 2007) Success :
No known problems apart from intermittent rain

Effelsberg (A. Polatidis at Mon Nov 5 10:49:57 2007) Success :

Medicina (Andrea Orlati at Mon Nov 5 12:41:53 2007) Success :
No known problems